OpenStudy (anonymous):

Algebra 2, Kinda easy but tricky... -5y - 9 = y - 1 ... Now what?

OpenStudy (anonymous):

Umm , 2 accounts? , same thing , take y to same side and -9 to other

OpenStudy (anonymous):

Possible answers: 1) -1/2 2) -2 1/2 3) -2 4) -2/5 and ima do that real quick

OpenStudy (anonymous):

Oh yeah, I did that and i got stuck, cause there's two ways u can go.

OpenStudy (anonymous):

-6y = 8 y = -8/6 = -4/3

OpenStudy (anonymous):

1) -5y - 9 = -y - 1 2) -5y - y = -9 - 1 3) -5y - y = -9 - 1 -y - y ------------------- -6y = -9 - 1 4) -6y = -9 - 1 -1 - 1 --------------- -6y = -10 Nvm bout that two options but it's not one of my possible answers. Possible answers: 1) -1/2 2) -2 1/2 3) -2 4) -2/5